# Streaming Federated Learning with Markovian Data

### v1

@@ Abstract
Federated learning (FL) is now recognized as a key framework for communication-efficient collaborative learning. Most theoretical and empirical studies, however, rely on the assumption that clients have access to pre-collected data sets, with limited investigation into scenarios where clients continuously collect data. In many real-world applications, particularly when data is generated by physical or biological processes, client data streams are often modeled by non-stationary Markov processes. Unlike standard i.i.d. sampling, the performance of FL with Markovian data streams remains poorly understood due to the statistical dependencies between client samples over time. In this paper, we investigate whether FL can still support collaborative learning with Markovian data streams. Specifically, we analyze the performance of Minibatch SGD, Local SGD, and a variant of Local SGD with momentum. We answer affirmatively under standard assumptions and smooth non-convex client objectives: the sample complexity is proportional to the inverse of the number of clients with a communication complexity comparable to the i.i.d. scenario. However, the sample complexity for Markovian data streams remains higher than for i.i.d. sampling.
@@ 1 Introduction
As edge networks grow in scale, clients are increasingly diverse [Lim et al., 2020]. Many of these clients are capable of continuously collecting and learning from data, instead of simply storing large quantities of historical data [Khodadadian et al., 2022]. This streaming data arises in applications ranging from health [Greco et al., 2019, Guo et al., 2024] and environmental monitoring [Hill and Minsker, 2010], to control of robots [Kober et al., 2013].
Each data stream provides detailed information about individual clients. However, applications such as environmental resource management [Sirimewan et al., 2024] or public health monitoring [Chae et al., 2018] often aim to derive population-level insights, leveraging data streams from multiple individuals to learn a global model. As this data may be sensitive or costly to communicate, centralized training solutions are undesirable. Federated Learning (FL) has emerged as a key strategy for communication-efficient collaborative learning [Kairouz et al., 2019]. In conventional FL scenarios, clients have access to local datasets used to train models locally. These models are then transmitted to a central server, where aggregation facilitates collaborative learning. FL minimizes communication overhead while learning models with high accuracy and ensuring client data are not shared.
Clients with streaming data must also cope with memory constraints [Marfoq et al., 2023, Liu et al., 2023]. This means that clients have limited control over data sampling: local updates can only be computed using the current data samples stored in memory. Moreover, data in health and environmental monitoring applications is often generated by non-linear physical and biological systems. The data generated by these systems is frequently modeled as non-stationary Markovian data streams [Anderson and Kurtz, 2015, Van Kampen, 1992]. When client memory caches are continually refreshed by Markovian streaming data, the samples available to compute updates are statistically dependent and non-stationary in time.
In this paper, we address the following question: can FL with Markovian data streams support collaborative learning? While Markovian sampling has been explored in the centralized setting [Even, 2023], existing work in the federated setting has focused on Federated Reinforcement Learning [Qi et al., 2021b, Khodadadian et al., 2022, Mangold et al., 2024] with convex objectives. However, classification and regression tasks in systems with non-linear dynamics often rely on deep neural networks [Yuan et al., 2020, Lucia et al., 2020], resulting in non-convex minimization problems during training. We therefore consider a family of training tasks beyond the work in [Qi et al., 2021b, Khodadadian et al., 2022, Mangold et al., 2024] characterized by $L$-smooth non-convex loss functions. Our main contributions can be summarized as follows:
@@
    Mitigating client heterogeneity: We show that a momentum-based variant of Local SGD introduced in the i.i.d. setting [Cheng et al., 2024, Yan et al., 2025] can effectively mitigate the client drift with Markovian data streams. In particular, unlike local SGD, local SGD with momentum matches the performance of Minibatch SGD up to a constant without requiring any bound on client data heterogeneity.
@@ 2 Related Works
Edge devices often encounter the challenge of processing continuous data streams while operating under memory constraints [Marfoq et al., 2023, Liu et al., 2023]. Learning from streaming data presents numerous challenges with solutions dependent on the data-generating process. One line of research addresses scenarios where the data distribution evolves over time, with continual learning methods designed to adapt the model to the changing data distribution [Schwarz et al., 2018]. Another line focuses on adversarial settings, where the data stream is generated by an adversary [Shalev-Shwartz, 2012], often addressed using online learning algorithms with regret guarantees.
@@
In this paper, we focus on learning from data generated by an unknown Markovian process. Markovian processes have been successfully applied to model various physical and biological systems [Anderson and Kurtz, 2015, Van Kampen, 1992]. Markovian data streams encompass the special case where data samples are drawn i.i.d. from an unknown distribution.
@@
In the centralized setting, this learning problem can be framed as a stochastic approximation problem, typically addressed by variants of stochastic gradient descent (SGD). While SGD with i.i.d. data is well-understood [Robbins and Monro, 1951, Lan, 2020], the Markovian nature of the data stream introduces challenges due to the temporal correlation between samples.
@@
For uniformly ergodic Markov chains, Duchi et al. [2011] first demonstrated that mirror descent achieves optimal convergence rates for Lipschitz, general convex, and non-convex problems. Recently, Dorfman and Levy [2022] proposed a random batch size algorithm that adapts to the mixing time of the Markov chain, while Beznosikov et al. [2023] adopted the same technique to develop accelerated methods and established lower bounds for strongly convex objectives.
@@
For general Markov chains, existing analysis of SGD-type algorithms [Sun et al., 2018, Doan et al., 2020b, a] have shown sub-optimal dependence on the mixing time and variance of stochastic gradient estimates relative to the lower bounds in Beznosikov et al. [2023], Duchi et al. [2011]. We refer to Beznosikov et al. [2023], Table 1 for an exhaustive review of the literature on SGD with Markovian noise.
@@
In the case of SGD with Markovian noise on a finite state space—a common scenario in distributed optimization [Bjorn Johansson, 2007]Even [2023], Hendrikx [2023] developed convergence theories and variance reduction techniques for first-order methods, demonstrating improvements over the gossip algorithm [Boyd et al., 2006].
@@
In the context of FL, a few papers have considered learning from data streams. Odeyomi and Zaruba [2021] and Yoon et al. [2021] framed the problem in an adversarial and in a continual learning setting, respectively. Chen et al. [2020] imposed strong homogeneity assumptions, such as requiring all clients to share a common optimal model. Marfoq et al. [2023] assumed data heterogeneity across clients but i.i.d. data streams at each client and considered the effect of heterogeous memory constraints. Liu et al. [2023] studied how to jointly tune batch sizes and number of local updates, but do not have convergence results in the streaming setting.
@@
In the case of Markovian data streams, recent work has primarily focused on Federated Reinforcement Learning (FRL) with linear function approximation. A key challenge is demonstrating that collaboration is beneficial in this setting; namely, that the per-client sample complexity of FRL algorithms decreases inversely with the number of clients. While this linear speed-up is well-understood for i.i.d. data, the impact of collaboration remains an open question in the presence of statistical dependence, as is the case with Markovian data streams. Khodadadian et al. [2022], Fabbro et al. [2023] demonstrated the benefits of cooperation under a strong homogeneity assumption on the client environment. Heterogeneous settings have been considered in Doan et al. [2019], Jin et al. [2022] with a linear speed-up established in Wang et al. [2024] subject to heterogeneity assumptions, which were then relaxed in Mangold et al. [2024] by incorporating control variates [Karimireddy et al., 2020].
@@
Beyond Markovian data streams, temporal correlations can arise from other factors in the FL setting. For instance, memory update strategies, as explored in Marfoq et al. [2023], can introduce dependencies between subsequent updates. Similarly, the availability of clients can exhibit Markovian behavior, as discussed in Ribero et al. [2023], Rodio et al. [2024], Sun et al. [2025]. While these sources of correlation are distinct from the Markovian data stream focus of this paper, they highlight the broader significance of understanding and mitigating temporal dependencies in the FL context.
@@ 3.1 Streaming Federated Learning
$F_{m}(w):=\mathbb{E}_{x\sim\pi_{m}}[f_{m}(w;x)],$ (1)
where $f_{m}:\mathbb{R}^{d}\times\Omega_{m}\rightarrow\mathbb{R}$ is a smooth function on $\mathbb{R}^{d}$ and $\pi_{m}$ is the target data distribution for client $m$. In the context of supervised learning, the objective $F_{m}$ corresponds to the true risk of the loss function $f_{m}$, parameterized by the parameter $w$, on a data sample $x$ drawn from $\pi_{m}$.
@@
$\min_{w\in\mathbb{R}^{d}}\frac{1}{M}\sum_{m=1}^{M}F_{m}(w)$ (2)
by iterating a two-phase procedure over multiple communication rounds. Two popular examples of this procedure are Minibatch SGD (Algorithm 2 in the Appendix) and Local SGD or FedAvg (Algorithm 3 in the Appendix) [Kairouz et al., 2019]. In the first phase of the $t$-th communication round, client $m$ has access to $K$ data samples $(x_{t}^{(m,k)})_{k\in[K]}$, which are used to compute a local update, corresponding to the gradient estimate $g_{t}^{m}$ in Minibatch SGD or iterate $w_{t}^{(m,K)}$ in Local SGD. In the second phase, the updates are aggregated via averaging.
In the standard FL scenario, the samples $(x_{t}^{(m,k)})_{k\in[K]}$ available at client $m$ in any communication round $t$ are fixed, corresponding to pre-collected data [Kairouz et al., 2019]. In contrast, in streaming FL, the data $(x_{t}^{(m,k)})_{k\in[K]}$ available to client $m$ can change in every communication round. Streaming FL models scenarios where clients have a limited memory cache of $K$ samples and new data is collected over time [Marfoq et al., 2023, Liu et al., 2023].
Streaming FL raises a key challenge: the data sampling process is not completely controlled by the clients. As a consequence, i.i.d. data samples from the target distribution $\pi_{m}$ may not be available. For example, patient data in health monitoring is continuously collected from wearable devices. Statistical dependence between data collected over time naturally arises due to latent factors such as daily routines, sleep patterns, and weather conditions [Li et al., 2017]. Moreover, the initial data collection or warm-up period [Baghdadi et al., 2021] might capture atypical behavior, such as adjustments to new wearable devices. As a consequence, the data stream of a patient $m$ will in general be non-stationary and not reflect the long-term target distribution $\pi_{m}$.
@@ 3.2 Markovian Data Streams
Data streams arising in health and environmental monitoring often exhibit specific statistical structures. Indeed, many biological, chemical, and physical systems are governed by non-stationary Markov processes [Anderson and Kurtz, 2015, Van Kampen, 1992]. Data streams arising in federated reinforcement learning also form Markov processes [Khodadadian et al., 2022, Qi et al., 2021a]; however, the use of stationary exploration policies leads to stationary data streams. stationary Markovian data stream.
Let $x_{t}^{m}=(x_{t}^{(m,k)})_{k\in[K]}$ be the data samples available to client $m$ in communication round $t$. We model the data stream $X_{m}=(x_{t}^{m},\leavevmode\nobreak\ t\in\mathbb{N})$ of client $m$ by a time-homogeneous Markov chain evolving on the state space $\Omega_{m}\subseteq\mathbb{R}^{d}$ with the corresponding Borel $\sigma$-field $\mathcal{B}_{m}$ Roberts and Rosenthal [2004], Meyn et al. [2009]. The evolution of the time-homogeneous Markov chain for client $m$ is characterized by the initial distribution $x_{0}^{m}\sim Q_{m}$ and by the transition kernel $P_{m}(x_{t}^{(m,k)},\mathrm{d}x_{t}^{(m,k+1)})$. The probability of transitioning from the state $x_{t}^{(m,k)}\in\Omega_{m}$ to a state $x_{t}^{(m,k+1)}\in\mathcal{X}_{m}\in\mathcal{B}_{m}$ is given by $\int_{x\in\mathcal{X}_{m}}P_{m}(x_{t}^{(m,k)},\mathrm{d}x)$. The $t$-step transition kernel for client $m$ from an initial state $x_{0}^{m}$ is denoted by $P_{m}^{t}(x_{0}^{m},\mathrm{d}x)$.
We focus on the case of independent clients, where $X_{m}$ is independent of $X_{m^{\prime}}$ for $m\neq m^{\prime}$. This scenario occurs when client data streams are generated by non-interacting processes. Key examples include patient health monitoring and environmental monitoring across spatially separated regions, where observations from different clients are often assumed to be independent [Sainani, 2010]. [^1] In the case of nearby or overlapping regions, statistical dependence between the data samples may arise Montero et al. [2015].
@@
$\int_{x\in\Omega_{m}}\pi_{m}(\mathrm{d}x)P_{m}(x,\mathrm{d}x_{t}^{(m,k)})=\pi_{m}(\mathrm{d}x_{t}^{(m,k)}).$ (3)
We say that $X_{m}$ is stationary if $x_{0}^{m}\sim\pi_{m}$; otherwise, $X_{m}$ is non-stationary.
The stationary distributions $\pi_{m},\leavevmode\nobreak\ m\in[M]$ correspond to the target distributions in (2), which capture the long-term statistics of the data samples. Since long-term statistics, rather than transient behavior of the samples, are of interest for learning in health and environmental monitoring scenarios, the stationary distribution is a natural choice for the target distribution. In order to guarantee that the Markov data stream associated with each client admits a stationary distribution, we impose the following assumption.
@@
The data samples of client $m\in[M]$ are drawn from an independent time-homogeneous Markov chain $X_{m}$ defined on $\left(\Omega_{m},\mathcal{B}_{m}\right)$ with transition kernel $P_{m}$ and initial distribution $\mathcal{Q}_{m}$, converging to the unique stationary distribution $\pi_{m}$.
Assumption 3.1 ensures that samples $x_{t}^{m}$ from the data stream of client $m$ will be approximately drawn from $\pi_{m}$ as $t$ diverges. As each client data stream evolves independently, the system-level data stream $X=((x_{t}^{m})_{m\in[M]},\leavevmode\nobreak\ t\in\mathbb{N})$ will eventually draw samples from the stationary distribution $\pi$ defined on $(\Omega:=\bigtimes_{m=1}^{M}\Omega_{m},\mathcal{B}:=\bigotimes_{m=1}^{M}\mathcal{B}_{m})$.
Let $Q:=\bigotimes_{m=1}^{M}Q_{m}$, $P(x_{0},\mathrm{d}x):=\prod_{m=1}^{M}P_{m}(x_{0}^{m},\mathrm{d}x^{m})$, and denote the $t$-step transition kernel for the system-level Markov chain initially in state $x_{0}$ by $P^{t}(x_{0},\mathrm{d}x)$. In the case $Q\neq\pi$, the system-level data stream is non-stationary. The distance of the distribution of $(x_{t}^{m})_{m\in[M]}$ and $\pi$ can be quantified by the mixing time of the system-level data stream $X$ in terms of the total variation norm.
Definition 3.2.
For two probability measures $\mu,\nu$ defined on $\left(\Omega,\mathcal{B}\right)$, the total variation distance between $\mu$ and $\nu$ is defined as:
$\left\lVert\mu-\nu\right\rVert_{TV}:=2\underset{A\in\mathcal{B}}{\sup}\left\lvert\mu(A)-\nu(A)\right\rvert.$ (4)
Definition 3.3.
For the Markov chain defined on $(\Omega,\mathcal{B})$ with transition kernel $P$ and unique stationary distribution $\pi$, let $d(t):=\underset{x\in\Omega}{\sup}\|P^{t}(x,\cdot)-\pi(\cdot)\|_{\mathrm{TV}}$. Then, the mixing time of the chain is defined as:
$\tau(\epsilon)$ $:=\inf\{t\geq 1:d(t)\leq\epsilon\}.$ (5)
We denote $\tau(1/4)$ simply as $\tau$ and, following convention, we call it the mixing time of the Markov chain.
Following prior work on stochastic optimization with Markovian data [Beznosikov et al., 2023, Nagaraj et al., 2020, Even, 2023], we assume that the system-level Markov chain is uniformly ergodic. This assumption ensures that the system-level Markov data stream admits exponential mixing, i.e., regardless of the initial distribution, it converges exponentially fast to the stationary distribution $\pi$.
@@
The system-level Markov chain $X$ has a finite mixing time $\tau$. Moreover, it is uniformly ergodic, that is, for every initial distribution $Q$ and $t\in\mathbb{N}$:
$\begin{split}\underset{(x_{1},\dots,x_{M})\in\Omega}{\sup}&\left\lVert QP^{t}((x_{1},\dots,x_{M}),.)-\pi(.)\right\rVert_{\mathrm{TV}}\\ &\leq\left(\frac{1}{4}\right)^{\lfloor t/\tau\rfloor}.\end{split}$ (6)

### v2

@@ Abstract
Federated learning (FL) is now recognized as a key framework for communication-efficient collaborative learning. Most theoretical and empirical studies, however, rely on the assumption that clients have access to pre-collected data sets, with limited investigation into scenarios where clients continuously collect data. In many real-world applications, particularly when data is generated by physical or biological processes, client data streams are often modeled by non-stationary Markov processes. Unlike standard i.i.d. sampling, the performance of FL with Markovian data streams remains poorly understood due to the statistical dependencies between client samples over time. In this paper, we investigate whether FL can still support collaborative learning with Markovian data streams. Specifically, we analyze the performance of Minibatch SGD, Local SGD, and a variant of Local SGD with momentum. We answer affirmatively under standard assumptions and smooth non-convex client objectives: the sample complexity is proportional to the inverse of the number of clients, with a communication complexity comparable to the i.i.d. scenario. However, the sample complexity for Markovian data streams remains higher than for i.i.d. sampling. Our analysis is validated via experiments with real pollution monitoring time series data.
@@ 1 Introduction
As edge networks grow in scale, clients are increasingly diverse [34]. Many of these clients are capable of continuously collecting and learning from data, instead of simply storing large quantities of historical data [28]. This streaming data arises in applications ranging from health [21, 22] and environmental monitoring [24], to control of robots [29].
Each data stream provides detailed information about individual clients. However, applications such as environmental resource management [54] or public health monitoring [6] often aim to derive population-level insights, leveraging data streams from multiple individuals to learn a global model. As this data may be sensitive or costly to communicate, centralized training solutions are undesirable. Federated Learning (FL) has emerged as a key strategy for communication-efficient collaborative learning [26]. In conventional FL scenarios, clients have access to local datasets used to train models locally. These models are then transmitted to a central server, where aggregation facilitates collaborative learning. FL minimizes communication overhead while learning models with high accuracy and ensuring client data are not shared.
Clients with streaming data must also cope with memory constraints [38, 35]. This means that clients have limited control over data sampling: local updates can only be computed using the current data samples stored in memory. Moreover, data in health and environmental monitoring applications is often generated by non-linear physical and biological systems. The data generated by these systems is frequently modeled as non-stationary Markovian data streams [1, 58]. When client memory caches are continually refreshed by Markovian streaming data, the samples available to compute updates are statistically dependent and non-stationary in time.
In this paper, we address the following question: Can FL with Markovian data streams support collaborative learning? While Markovian sampling has been explored in the centralized setting [17], existing work in the federated setting has focused on Federated Reinforcement Learning [45, 28, 37] with convex objectives. However, classification and regression tasks in systems with non-linear dynamics often rely on deep neural networks or non-convex regularization [63, 36], resulting in non-convex minimization problems during training. We therefore consider a family of training tasks beyond the work in [45, 28, 37] characterized by smooth non-convex loss functions. Our main contributions can be summarized as follows:
@@
    Mitigating client heterogeneity: We show that a momentum-based variant of Local SGD introduced in the i.i.d. setting [9, 61] can effectively mitigate the client drift with Markovian data streams. In particular, unlike Local SGD, Local SGD with Momentum matches the performance of Minibatch SGD up to a constant without requiring any bound on client data heterogeneity.
  3. (iii)
    Validation on Environmental Monitoring Data: The key difference between i.i.d. and Markovian sampling is statistical dependence between consecutive samples. We validate our analysis using multi-site pollution time series data [7], which shows the benefit of collaboration on real data with dependent samples.
@@ 2 Related Works
Edge devices often encounter the challenge of processing continuous data streams while operating under memory constraints [38, 35]. Learning from streaming data presents numerous challenges, with solutions dependent on the data-generating process. One line of research addresses scenarios where the data distribution evolves over time, with continual learning methods designed to adapt the model to the changing data distribution [51]. Another line focuses on adversarial settings, where the data stream is generated by an adversary [53], often addressed using online learning algorithms with regret guarantees.
@@
In this paper, we focus on learning from data generated by Markovian processes. Markovian processes have been successfully applied to model various physical and biological systems [1, 58], and encompass the special case where data samples are drawn i.i.d. from an unknown distribution.
@@
In the centralized setting, this learning problem can be framed as a stochastic approximation problem, typically addressed by variants of stochastic gradient descent (SGD). While SGD with i.i.d. data is well-understood [47, 31], the Markovian nature of the data stream introduces challenges due to the temporal correlation between samples.
@@
For uniformly ergodic Markov chains, [16] first demonstrated that mirror descent achieves optimal convergence rates for Lipschitz, general convex, and non-convex problems. Recently, [15] proposed a random batch size algorithm that adapts to the mixing time of the Markov chain, while [3] adopted the same technique to develop accelerated methods and established lower bounds for strongly convex objectives.
@@
For general Markov chains, existing analysis of SGD-type algorithms [55, 14, 13] have shown sub-optimal dependence on the mixing time and on the variance of stochastic gradient estimates relative to the lower bounds in [3, 16]. We refer to [3, Table 1] for an exhaustive review of the literature on SGD with Markovian noise.
@@
In the case of SGD with Markovian noise on a finite state space—a common scenario in distributed optimization [4][17, 23] developed convergence theories and variance reduction techniques for first-order methods, demonstrating improvements over the gossip algorithm [5].
@@
In the context of FL, a few papers have considered learning from data streams. [41] and [62] framed the problem in an adversarial and in a continual learning setting, respectively. [8] imposed strong homogeneity assumptions, such as requiring all clients to share a common optimal model. [38] assumed that each client’s data stream consists of i.i.d. samples, while allowing for heterogeneous data distributions across clients. The study focused on temporal dependencies arising from the use of local memory and analyzed the impact of heterogeneous memory constraints. [35] studied how to jointly tune batch sizes and number of local updates, but did not have convergence results in the streaming setting.
@@
In the case of Markovian data streams, recent work has primarily focused on Federated Reinforcement Learning (FRL) with linear function approximation. A key challenge is demonstrating that collaboration is beneficial in this setting; namely, that the per-client sample complexity of FRL algorithms decreases inversely with the number of clients. While this linear speed-up is well-understood for i.i.d. data, the impact of collaboration remains an open question in the presence of statistical dependence, as is the case with Markovian data streams. [28, 18] demonstrated the benefits of cooperation under a strong homogeneity assumption on the client environment. Heterogeneous settings have been considered in [12, 25] with a linear speed-up established in [59] subject to heterogeneity assumptions, which were then relaxed in [37] by incorporating control variates [27]. Unfortunately, these results do not readily extend beyond the FRL setting.
@@
Beyond Markovian data streams, temporal correlations can arise from other factors in the FL setting. For instance, memory update strategies, as explored in [38], can introduce dependencies between subsequent updates. Similarly, the availability of clients can exhibit Markovian behavior, as discussed in [46, 49, 56]. While these sources of correlation are distinct from the Markovian data stream focus of this paper, they highlight the broader significance of understanding and mitigating temporal dependencies in the FL context.
@@ 3 Problem Setup
In this section, we introduce our problem setting. We denote by $M]\coloneqq[1,M]$ the set of positive integers up to $M$, and by $K]_{0}\coloneqq[0,K-1]$ the set of non-negative integers less than $K$.
@@ 3.1 Streaming Federated Learning
$F_{m}(w)=\mathbb{E}_{x\sim\pi_{m}}[f_{m}(w;x)],$ (1)
where $f_{m}:\mathbb{R}^{d}\times\Omega_{m}\rightarrow\mathbb{R}$ and $\pi_{m}$ is the target data distribution. for client $m$. In the context of supervised learning, the objective $F_{m}$ corresponds to the true risk of the loss function $f_{m}$, parameterized by the parameter $w$, on a data sample $x$ drawn from $\pi_{m}$.
@@
$\min_{w\in\mathbb{R}^{d}}\frac{1}{M}\sum_{m=1}^{M}F_{m}(w),$ (2)
by iterating a two-phase procedure over multiple communication rounds. Two popular examples of this procedure are Minibatch SGD (Algorithm 2 in Appendix A) and Local SGD or FedAvg (Algorithm 3 in Appendix A) [26]. In the first phase of the $t$-th communication round, client $m$ has access to $K$ data samples. These samples are used to compute a local update: either a gradient estimate $g_{t}^{(m)}$ in the case of Minibatch SGD, or a model iterate $w_{t}^{(m,K)}$ obtained after performing $K$ local gradient steps in the case of Local SGD. In the second phase, the server aggregates the local updates from all clients—typically via averaging—to update the global model. In the second phase, the updates are aggregated via averaging.
In the standard FL scenario, the samples available at client $m$ in any communication round $t$ are fixed, corresponding to pre-collected data [26]. In contrast, in streaming FL, the data available to client $m$ can change in every communication round. Streaming FL models scenarios where clients have a limited memory cache of $K$ samples and new data is collected over time [38, 35].
Streaming FL raises a key challenge: the data sampling process is not completely controlled by the clients. As a consequence, i.i.d. data samples from the target distribution $\pi_{m}$ may not be available. For example, patient data in health monitoring is continuously collected from wearable devices. Statistical dependence between data collected over time naturally arises due to latent factors such as daily routines, sleep patterns, and weather conditions [32]. Moreover, the initial data collection or warm-up period [2] might capture atypical behavior, such as adjustments to new wearable devices. As a consequence, the data stream of a patient $m$ will, in general, be non-stationary and not reflect the long-term target distribution $\pi_{m}$.
@@ 3.2 Markovian Data Streams
Data streams arising in health and environmental monitoring often exhibit specific statistical structures. Indeed, many biological, chemical, and physical systems are governed by non-stationary Markov processes [1, 58]. Data streams arising in federated reinforcement learning also form Markov processes [28, 44]; however, the use of stationary exploration policies leads to stationary data streams.
We model the data stream $X_{m}=\left(x^{m}_{t}\right)_{t\in\mathbb{N}}$ of client $m$ by a time-homogeneous Markov chain evolving on the state space $\Omega_{m}\subseteq\mathbb{R}^{d}$ with the corresponding Borel $\sigma$-field $\mathcal{B}_{m}$ [48, 39]. As an abuse of notation, for all $t\in\mathbb{N}$, and for all $k\in[K]_{0}$, we denote by $x_{t}^{(m,k)}\coloneqq x^{m}_{Kt+k}$ the $k$-th sample available to client $m$ at the $t$-th communication round. We may also use $x_{t}^{m}\coloneqq\left(x_{t}^{(m,k)}\right)_{k\in[K]_{0}}$ to denote all $K$ data samples available to client $m$ in communication round $t$. The evolution of the time-homogeneous Markov chain for client $m$ is characterized by the initial distribution $x_{0}^{(m,0)}\sim\mu_{m}$ and by the transition kernel $P_{m}$.
We focus on the case of independent clients, where $X_{m}$ is independent of $X_{m^{\prime}}$ for $m\neq m^{\prime}$. This scenario occurs when client data streams are generated by non-interacting processes. Key examples include patient health monitoring and environmental monitoring across spatially separated regions, where observations from different clients are often assumed to be independent [50]. [^1] In the case of nearby or overlapping regions, statistical dependence between the data samples may arise [40].
@@
$\int_{x\in\Omega_{m}}\mathrm{d}\pi_{m}(x)P_{m}(x,\mathrm{d}x_{t}^{(m,k)})=\mathrm{d}\pi_{m}(x_{t}^{(m,k)}).$ (3)
We say that $X_{m}$ is stationary if $x_{0}^{(m,0)}\sim\pi_{m}$; otherwise, $X_{m}$ is non-stationary.
The stationary distributions $\pi_{m},~m\in[M]$ correspond to the target distributions in (2), which capture the long-term statistics of the data samples. Since long-term statistics, rather than transient behavior of the samples, are of interest for learning in health and environmental monitoring scenarios, the stationary distribution is a natural choice for the target distribution. Hence, we make the following assumption, which guarantees that the Markov process associated with each client admits a unique stationary distribution.
@@
The data samples of client $m\in[M]$ are drawn from an independent time-homogeneous Markov chain $X_{m}$ defined on $\left(\Omega_{m},\mathcal{B}_{m}\right)$ with transition kernel $P_{m}$ and initial distribution $\mu_{m}$, converging to the unique stationary distribution $\pi_{m}$.
Assumption 3.1 ensures that samples $x_{t}^{m}$ from the data stream of client $m$ will be approximately drawn from $\pi_{m}$ as $t$ diverges.
We also define by $X=\left(X_{m}\right)_{m\in[M]}$ the system-level Markov process defined on $\left(\Omega,\mathcal{B}\right)$, where $\Omega\coloneqq\bigtimes_{m=1}^{M}\Omega_{m}$ and $\mathcal{B}\coloneqq\bigotimes_{m=1}^{M}\mathcal{B}_{m}$. This Markov chain, at each time step, evolves independently on each of $M$ coordinates according to the corresponding transition kernel $P_{m}$. We denote by $P$ its transition matrix. Similarly, for the ease of notation, we write $x_{t}^{k}\coloneqq\left(x_{t}^{(m,k)}\right)_{m\in[M]}$ and $x_{t}\coloneqq\left(x_{t}^{k}\right)_{k\in[K]_{0}}$.
In Proposition B.6, we provide a formal characterization of $P$ and we also prove that $X$ is indeed a well-defined Markov chain. Furthermore, under Assumption 3.1, $X$ admits a stationary distribution $\pi=\bigotimes_{m=1}^{M}\pi_{m}$. We denote by $\nu_{ps}$ the pseudo spectral gap of $P$ [43, Section 3.1].
We also make the following assumption on the absolute continuity of the transition kernel $P$ with respect to the stationary distribution $\pi$.
@@
For every $x\in\Omega$, the probability measure $P(x,.)$ is absolutely continuous with respect to the stationary measure $\pi$, and its Radon-Nikodym derivative $\frac{dP(x,.)}{d\pi}$ is uniformly bounded on sets with non-zero measure with respect to $\pi$. In particular,
$C_{\infty}=\underset{x\in\Omega}{\sup}\operatorname*{ess\,sup}\left\lvert\frac{dP(x,.)}{d\pi}\right\rvert$$\infty$
We note that Assumption 3.2 is generally mild, in the sense that it only requires, for every $x\in\Omega$, $P(x,.)$ is absolutely continuous with respect to $\pi$, and every $\pi$-integrable function is also $P(x,.)$-integrable. This assumption is commonly used in the analysis of other stochastic approximation algorithms with Markovian noise, for example, Markov chain Monte Carlo [19, Theorem 12]. In the case where $\Omega$ is finite, $C_{\infty}=\max_{x,y\in\Omega}\frac{P(x,y)}{\pi(y)}$. We provide further discussion of the dependence of $C_{\infty}$ on the transition kernel in Appendix D.5.
In contrast to most existing studies on stochastic optimization (approximation) with Markovian data, we do not impose any assumption on the speed of convergence to the stationary measure of the underlying Markov processes, such as the commonly used uniform geometric ergodicity assumption (see, e.g., 3, Assumption A3; 37, Assumption A3; 59, Assumption 3; 65, Assumption 3), which requires exponentially fast convergence to the stationary measure. Prior works rely on this fast mixing property to handle the non-stationarity of Markovian data when taking conditional expectations, which is ubiquitous in the analysis of SGD-type algorithms. However, our analysis shows that, under the mild Assumption 3.2, a simple change of measures suffices. As uniform geometric ergodicity is typically difficult to verify in real-world scenarios, our results are therefore more broadly applicable. For completeness, in Appendix D.4, we provide an extension of our analysis under the stronger uniform geometric ergodicity assumption.